# Latent space models for grouped multiplex networks

### v1

@@ B.4 Technical lemmas
Lemma B.1.
Let $E_{k\ell}\in\mathbb{R}^{n\times n},(k,\ell)\in\mathcal{I}$ be symmetric Gaussian matrices with independent entries following
$E_{k\ell,ij}\stackrel{{\scriptstyle\text{{ind}}}}{{\sim}}\mathcal{N}(0,\sigma^{2}),\qquad 1\leq i\leq j\leq n,\qquad(k,\ell)\in\mathcal{I}.$
Then, the following concentration bounds hold simultaneously for the operator norms of the individual matrices $E_{k\ell},\ (k,\ell)\in\mathcal{I},$ their group-wise averages $\bar{E}_{k}={1\over m_{k}}\sum_{\ell=1}^{m_{k}}E_{k\ell},\ k=1,\ldots,K$, and the grand-total average $\bar{E}={1\over M}\sum_{(k,\ell)\in\mathcal{I}}E_{k\ell}$:
$\mathbb{P}\Bigl(\|E_{k\ell}\|_{2}\leq 3\sigma\sqrt{n},\ (k,\ell)\in\mathcal{I};\ \|\bar{E}_{k}\|_{2}\leq 3\sigma\sqrt{n/m_{k}};\ \|\bar{E}\|_{2}\leq 3\sigma\sqrt{n/M}\Bigr)\geq 1-(M+K+1)ne^{-C_{0}n},$
where $C_{0}>0$ is some universal constant.
Proof.
Analogously to Lemma 3 in (multiness). ∎
Lemma B.2 (Version of Theorem 1, (cai_zhang)).
@@
$Z=\left[\begin{array}[]{ll}V&V_{\perp}\end{array}\right]\cdot\left[\begin{array}[]{cc}\Gamma_{1}&0\\ 0&\Gamma_{2}\end{array}\right]\cdot\left[\begin{array}[]{c}V^{\top}\\ V_{\perp}^{\top}\end{array}\right]$
and
$\tilde{Z}=\left[\begin{array}[]{ll}\tilde{V}&\tilde{V}_{\perp}\end{array}\right]\cdot\left[\begin{array}[]{cc}\tilde{\Gamma}_{1}&0\\ 0&\tilde{\Gamma}_{2}\end{array}\right]\cdot\left[\begin{array}[]{c}\tilde{V}^{\top}\\ \tilde{V}_{\perp}^{\top}\end{array}\right$],
where $V^{\top}V=I_{d},\ V_{\perp}^{\top}V_{\perp}=I_{n-d},\ \Gamma_{1}=\operatorname{diag}[\gamma_{1}(Z),\ldots,\gamma_{d}(Z)],\ \Gamma_{2}=\operatorname{diag}[\gamma_{d+1}(Z),\ldots,\gamma_{n}(Z)]$ for some $d<n$, and $\tilde{\Gamma}_{1},\tilde{\Gamma}_{2},\tilde{V},\tilde{V}_{\perp}$ defined similarly. Let also
@@
Condition (56) is stronger than the one in Theorem 1, (cai_zhang) since it implies
@@
and we can further loosen the original upper-bound on $\|\sin\Theta(V,\hat{V})\|_{2}$ as follows
$\|\sin\Theta(V,\hat{V})\|_{2}\leq$ $\ {\alpha e_{12}+\beta e_{21}\over\alpha^{2}-\beta^{2}-e_{12}^{2}\wedge e_{21}^{2}}$
$\leq$ $\ {(\alpha+\beta+e_{12}\wedge e_{21})(e_{12}\vee e_{21})\over\alpha^{2}-(\beta+e_{12}\wedge e_{21})^{2}}$
@@
With notations of Lemma B.2, if $d$ is the rank of $Z$ and $|\gamma_{d}(Z)|>3\|E\|_{2}$, then
$\|\sin\Theta(V,\tilde{V})\|_{2}\leq{\|E\|_{2}\over|\gamma_{d}(Z)|-3\|E\|_{2}}.$
@@
Notice that $e_{12},e_{21}\leq\|E\|_{2}$ and by Weyl’s inequality combined with submultiplicativity,
$\alpha=|\gamma_{\min}(V^{\top}\tilde{Z}V)|\geq|\gamma_{d}(V^{\top}ZV+V^{\top}EV)|\geq|\gamma_{d}(Z)|-\|E\|_{2}$
$\beta=\|V_{\perp}^{\top}\tilde{Z}V_{\perp}\|_{2}\leq\|V_{\perp}^{\top}ZV_{\perp}\|_{2}+\|V^{\top}EV\|_{2}\leq\|E\|_{2}.$
@@
The next lemma summarizes various relationships between a matrix and the soft-thresholding of its perturbation.
@@
    The spectral norm of their difference can be bounded as
@@
    If $\rho\geq\|E\|_{2}$, the Frobenius norm of their difference satisfies
@@
    If $\rho\geq\|E\|_{2}$ and $Z$ is positive semi-definite (PSD), then $\hat{Z}$ is also PSD.
@@
    If $\rho\geq\|E\|_{2}+2\|Z\|_{2}\|\sin\Theta(V,\tilde{V})\|_{2}^{2}$, the rank of $\hat{Z}$ is at most $d$.
@@
    Consider the eigendecomposition $Z+E=\tilde{V}\Sigma\tilde{V}^{\top}$. Then, the eigendecomposition of $\hat{Z}$ can be written as $\hat{Z}=\tilde{V}\Sigma_{\rho}\tilde{V}^{\top}$ with
    $\Sigma_{\rho}=\operatorname{diag}\left[\operatorname{sign}(\Sigma_{11})(|\Sigma_{11}|-\rho)_{+},\ldots,\operatorname{sign}(\Sigma_{nn})(|\Sigma_{nn}|-\rho)_{+}\right].$
@@
    $\hat{Z}-Z=\tilde{U}\Sigma_{\rho}\tilde{V}^{\top}-\left(\tilde{U}\Sigma\tilde{V}^{\top}-E\right)=\tilde{U}\left(\Sigma_{\rho}-\Sigma\right)\tilde{V}^{\top}+E$
    where $\Sigma_{\rho}-\Sigma$ is diagonal with
    $\Sigma_{\rho}-\Sigma]_{ii}=\operatorname{sign}(\Sigma_{ii})(|\Sigma_{ii}|-\rho)_{+}-\Sigma_{ii}=-\operatorname{sign}(\Sigma_{ii})\min(|\Sigma_{ii}|,\rho),$
@@
    $\|\hat{Z}-Z\|_{2}\leq\|\Sigma_{\rho}-\Sigma\|_{2}+\|E\|_{2}=\max_{1\leq i\leq n}|\min\{\rho,|\Sigma_{ii}|\}|+\|E\|_{2}\leq\rho+\|E\|_{2}.$
@@
    $\hat{Z}=\operatorname*{argmin}_{y\in\mathbb{R}^{n\times n}}\Bigl\{{1\over 2}\|Y-Z\|_{F}^{2}+\rho\|Y\|_{*}\Bigr\}$
@@
    $\|\hat{Z}\|_{*}-\|\check{Z}\|_{*}\geq\left\langle\check{G}_{Z},\hat{Z}-\check{Z}\right\rangle$
    $\|\check{Z}\|_{*}-\|\hat{Z}\|_{*}\geq\left\langle G_{Z},\check{Z}-\hat{Z}\right\rangle$
@@
    $\langle\hat{Z}-Z,\hat{Z}-\check{Z}\rangle\leq-\rho\langle\check{G}_{Z},\hat{Z}-\check{Z}\rangle+\langle E,\hat{Z}-\check{Z}\rangle.$ (61)
    which is the inequality of the same form as (58) but now, instead of a fixed $G_{Z}$, we have a free choice of $\check{G}_{Z}$. By Koltchinskii2010, subgradient $\check{G}_{Z}$ can be expressed as
    $\check{G}_{Z}=\sum_{i=1}^{d}u_{i}v_{i}^{\top}+\check{\mathcal{P}}_{Z}^{\perp}W\check{\mathcal{P}}_{Z}^{\perp}$
    where $\check{\mathcal{P}}_{Z}$ is the projector on the column space of $\check{Z}$, vectors $u_{i},v_{i}$ are left and right singular vectors of $\check{Z}$ respectively, and $W$ is an arbitrary matrix satisfying $\|W\|_{2}\leq 1$. So, we can specify $W$ to attain the upper bound of the following expression obtained using the duality of nuclear and operator norms
@@
    $|\langle\sum_{i=1}^{d}u_{i}v_{i}^{\top},\hat{Z}-\check{Z}\rangle|\leq\sqrt{d}\|\hat{Z}-\check{Z}\|_{F}$ (63)
@@
    $\leq 3\sqrt{d}\left\|E\right\|_{2}\|\hat{Z}-\check{Z}\|_{F}+\left\|E\right\|_{2}\|\mathcal{P}_{Z}^{\perp}\hat{Z}\mathcal{P}_{Z}^{\perp}\|_{*}.$
@@
    $\|\hat{Z}-Z\|_{F}^{2}$ $+\left(\rho-\left\|E\right\|_{2}\right)\|\check{\mathcal{P}}_{Z}^{\perp}\hat{Z}\check{\mathcal{P}}_{Z}^{\perp}\|_{*}\leq\sqrt{d}\left(\rho+3\left\|E\right\|_{2}\right)\|\hat{Z}-Z\|_{F}$
@@
    $\left|v^{\top}Zv\right|=\left|v^{\top}VV^{\top}ZVV^{\top}v\right|\leq\left\|Z\right\|_{2}\left\|V^{\top}v\right\|_{2}^{2}$ (65)
@@
    So, at most first $d$ eigenvalues of $\hat{Z}$ can survive the soft-thresholding step
@@
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}(\hat{Z}_{\ell}-Z_{\ell})\Bigr\|_{2}\leq 11(\gamma\theta\vee\rho)\Bigl[{1\over m}\vee s\vee\theta\Bigr]^{1/2}$ (67)
@@
Term $(\mathcal{I})$. We start by bounding the cosine similarity between perturbed eigenspaces in terms of the similarity of the true eigenspaces. By (42) and the triangular inequality, we have
$\left\|\tilde{V}_{\ell_{1}}^{\top}\tilde{V}_{\ell_{2}}\right\|_{2}$ $=\inf_{O_{1}\in\mathcal{O}_{d_{\ell_{1}}},\ O_{2}\in\mathcal{O}_{d_{\ell_{2}}}}\left\|\tilde{V}_{\ell_{1}}^{\top}\tilde{V}_{\ell_{2}}-\tilde{V}_{\ell_{1}}^{\top}V_{\ell_{2}}O_{1}+\tilde{V}_{\ell_{1}}^{\top}V_{\ell_{2}}O_{1}-O_{2}V_{\ell_{1}}^{\top}V_{\ell_{2}}O_{1}+O_{2}V_{\ell_{1}}^{\top}V_{\ell_{2}}O_{1}\right\|_{2}$ (68)
$\leq\inf_{O_{1}\in\mathcal{O}_{d_{\ell_{1}}}}\left\|\tilde{V}_{\ell_{2}}-V_{\ell_{2}}O_{1}\right\|_{2}+\inf_{O_{2}\in\mathcal{O}_{d_{\ell_{2}}}}\left\|\tilde{V}_{\ell_{1}}-V_{\ell_{1}}O_{2}\right\|_{2}+\left\|V_{\ell_{1}}^{\top}V_{\ell_{2}}\right\|_{2}$
$\leq 2\sqrt{2}\theta+s$
Notice that by property 3 in Lemma B.3 and the threshold choice in (66), we have $\|\hat{Z}_{\ell}\|_{2}\leq\gamma$. Thus, by (41) and submultiplicativity
$\left\|\left(\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}-V_{\ell}V_{\ell}^{\top}\right)\hat{Z}_{\ell}\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}\right\|_{2}$ $\leq\|\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}-V_{\ell}V_{\ell}^{\top}\|_{2}\|\hat{Z}_{\ell}\|_{2}\leq 2\theta\gamma$
@@
Therefore, by Lemma 4 in multiness
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}(\mathcal{I})_{\ell}\Bigr\|_{2}\leq 2m^{-1/2}\theta\gamma\Bigl[1+m(2\sqrt{2}\theta+s)\Bigr]^{1/2}\leq 5\theta\gamma\Bigl[{1\over m}\vee\theta\vee s\Bigr]^{1/2},$
where in the last line we used $(2+2\sqrt{2})^{1/2}<2.5$.
Term $(\mathcal{II})$. With the choice of the thresholds in (B.3), we have by property 1 in Lemma B.3 that $\|\Delta_{\ell}\|_{2}\leq 2\rho$
$\|V_{\ell_{1}}V_{\ell_{1}}^{\top}\Delta_{\ell_{1}}V_{\ell_{1}}V_{\ell_{1}}^{\top}\cdot V_{\ell_{2}}V_{\ell_{2}}^{\top}\Delta_{\ell_{2}}V_{\ell_{2}}V_{\ell_{2}}^{\top}\|_{2}\leq\|\Delta_{\ell_{1}}\|_{2}\|V_{\ell_{1}}^{\top}V_{\ell_{2}}\|_{2}\|\Delta_{\ell_{2}}\|_{2}\leq 4s\rho^{2}$
Thus, by Lemma 4 in multiness,
$\Bigl\|{1\over m}\sum_{k=1}^{m}(\mathcal{II})_{k}\Bigr\|_{2}\leq 2m^{-1/2}\rho\Bigl[1+ms\Bigr]^{1/2}\leq 3\rho\Bigl[{1\over m}\vee s\Bigr]^{1/2}$
Term $(\mathcal{III})$. Similarly to Term $(\mathcal{I})$, we can establish by substituting $\|\tilde{V}_{\ell_{1}}\tilde{V}_{\ell_{2}}\|_{2}$ with $\|V_{\ell_{1}}V_{\ell_{2}}\|_{2}$ in (69):
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}(\mathcal{III})_{\ell}\Bigr\|_{2}\leq 2m^{-1/2}\theta\gamma\Bigl[1+ms\Bigr]^{1/2}\leq 3\theta\gamma\Bigl[{1\over m}\vee s\Bigr]^{1/2}.$

### v2

@@ B.4 Technical lemmas
Lemma B.1 (Version of Theorem 1, (Cai and Zhang, 2016)).
@@
$Z=\left[\begin{array}[]{ll}V&V_{\perp}\end{array}\right]\left[\begin{array}[]{cc}\Gamma_{1}&0\\ 0&\Gamma_{2}\end{array}\right]\left[\begin{array}[]{c}V^{\top}\\ V_{\perp}^{\top}\end{array}\right],\ \ \ \tilde{Z}=\left[\begin{array}[]{ll}\tilde{V}&\tilde{V}_{\perp}\end{array}\right]\left[\begin{array}[]{cc}\tilde{\Gamma}_{1}&0\\ 0&\tilde{\Gamma}_{2}\end{array}\right]\left[\begin{array}[]{c}\tilde{V}^{\top}\\ \tilde{V}_{\perp}^{\top}\end{array}\right$],
where $V^{\top}V=I_{d},\ V_{\perp}^{\top}V_{\perp}=I_{n-d},\ \Gamma_{1}=\operatorname{diag}[\gamma_{1}(Z),\ldots,\gamma_{d}(Z)],\ \Gamma_{2}=\operatorname{diag}[\gamma_{d+1}(Z),\ldots,\gamma_{n}(Z)]$ for some $d<n$, and $\tilde{\Gamma}_{1},\tilde{\Gamma}_{2},\tilde{V},\tilde{V}_{\perp}$ are defined similarly. Let
@@
Condition (48) is stronger than the one in Theorem 1, (Cai and Zhang, 2016) since it implies
@@
and we can further relax the original upper bound on $\|\sin\Theta(V,\hat{V})\|_{2}$ as follows:
$\|\sin\Theta(V,\hat{V})\|_{2}$ $\leq\ {\alpha e_{12}+\beta e_{21}\over\alpha^{2}-\beta^{2}-e_{12}^{2}\wedge e_{21}^{2}}\leq\ {(\alpha+\beta+e_{12}\wedge e_{21})(e_{12}\vee e_{21})\over\alpha^{2}-(\beta+e_{12}\wedge e_{21})^{2}}$
@@
In notation of Lemma B.1, if $d$ is the rank of $Z$ and $|\gamma_{d}(Z)|>3\|E\|_{2}$, then
$\|\sin\Theta(V,\tilde{V})\|_{2}\leq{\|\mathcal{P}_{V}E\|_{2}\over|\gamma_{d}(Z)|-3\|E\|_{2}}.$
@@
By $e_{12},e_{21}\leq\|\mathcal{P}_{V}E\|_{2}$ and Weyl’s inequality combined with submultiplicativity,
$\alpha$ $=|\gamma_{\min}(V^{\top}\tilde{Z}V)|\geq|\gamma_{d}(V^{\top}ZV+V^{\top}EV)|\geq|\gamma_{d}(Z)|-\|E\|_{2},$
$\beta$ $=\|V_{\perp}^{\top}\tilde{Z}V_{\perp}\|_{2}\leq\|V_{\perp}^{\top}ZV_{\perp}\|_{2}+\|V_{\perp}^{\top}EV_{\perp}\|_{2}\leq\|E\|_{2}.$
@@
The next lemma lists various relationships between a matrix and the soft thresholding of its perturbation.
@@
    The spectral norm of the difference can be bounded as
@@
    If $\rho\geq\|E\|_{2}$, the Frobenius norm of the difference satisfies
@@
    If $\rho\geq\|E\|_{2}$ and $Z$ is PSD, then $\hat{Z}$ is also PSD.
@@
    If $\rho\geq\|E\|_{2}+2\|Z\|_{2}\|\sin\Theta(V,\tilde{V})\|_{2}^{2}$, then $\mathrm{rank}(\hat{Z})\leq d$.
@@
    Consider the eigendecomposition $\tilde{Z}=Z+E=\tilde{U}\tilde{\Gamma}\tilde{U}^{\top}$. Then, the eigendecomposition of $\hat{Z}$ can be written as $\hat{Z}=\tilde{U}\tilde{\Gamma}_{\rho}\tilde{U}^{\top}$ with
    $\tilde{\Gamma}_{\rho}=\operatorname{diag}\left[\operatorname{sign}(\tilde{\Gamma}_{11})(|\tilde{\Gamma}_{11}|-\rho)_{+},\ldots,\operatorname{sign}(\tilde{\Gamma}_{nn})(|\tilde{\Gamma}_{nn}|-\rho)_{+}\right].$
@@
    $\hat{Z}-Z=\tilde{U}\tilde{\Gamma}_{\rho}\tilde{U}^{\top}-\left(\tilde{U}\tilde{\Gamma}\tilde{U}^{\top}-E\right)=\tilde{U}\left(\tilde{\Gamma}_{\rho}-\tilde{\Gamma}\right)\tilde{U}^{\top}+E$
    where $\tilde{\Gamma}_{\rho}-\tilde{\Gamma}$ is diagonal with
    $\tilde{\Gamma}_{\rho}-\tilde{\Gamma}]_{ii}=\operatorname{sign}(\tilde{\Gamma}_{ii})(|\tilde{\Gamma}_{ii}|-\rho)_{+}-\tilde{\Gamma}_{ii}=-\operatorname{sign}(\tilde{\Gamma}_{ii})\min(|\tilde{\Gamma}_{ii}|,\rho),$
@@
    $\|\hat{Z}-Z\|_{2}\leq\|\tilde{\Gamma}_{\rho}-\tilde{\Gamma}\|_{2}+\|E\|_{2}=\max_{1\leq i\leq n}|\min\{\rho,|\tilde{\Gamma}_{ii}|\}|+\|E\|_{2}\leq\rho+\|E\|_{2}.$
@@
    $\hat{Z}=\operatorname*{argmin}_{Y\in\mathbb{R}^{n\times n}}\Bigl\{{1\over 2}\|Y-(Z+E)\|_{F}^{2}+\rho\|Y\|_{*}\Bigr\}$
@@
    $\|\hat{Z}\|_{*}-\|\check{Z}\|_{*}\geq\left\langle\check{G}_{Z},\hat{Z}-\check{Z}\right\rangle,\ \ \|\check{Z}\|_{*}-\|\hat{Z}\|_{*}\geq\left\langle G_{Z},\check{Z}-\hat{Z}\right\rangle$
@@
    $\langle\hat{Z}-Z,\hat{Z}-\check{Z}\rangle\leq-\rho\langle\check{G}_{Z},\hat{Z}-\check{Z}\rangle+\langle E,\hat{Z}-\check{Z}\rangle,$ (53)
    which is the inequality of the same form as (50) but now instead of a fixed $G_{Z}$ we have a free choice of $\check{G}_{Z}$. By Koltchinskii et al. (2010), subgradient $\check{G}_{Z}$ can be expressed as
    $\check{G}_{Z}=\sum_{i=1}^{d}u_{i}v_{i}^{\top}+\check{\mathcal{P}}_{Z}^{\perp}W\check{\mathcal{P}}_{Z}^{\perp},$
    where $\check{\mathcal{P}}_{Z}$ is the projector on the column space of $\check{Z}$, vectors $u_{i},v_{i}$ are left and right singular vectors of $\check{Z}$, respectively, and $W$ is an arbitrary matrix satisfying $\|W\|_{2}\leq 1$. We can specify $W$ to attain the upper bound of the following expression obtained using the duality of nuclear and operator norms
@@
    $|\langle\sum_{i=1}^{d}u_{i}v_{i}^{\top},\hat{Z}-\check{Z}\rangle|\leq\sqrt{d}\|\hat{Z}-\check{Z}\|_{F}.$ (55)
@@
    $\leq 3\sqrt{d}\left\|E\right\|_{2}\|\hat{Z}-\check{Z}\|_{F}+\left\|E\right\|_{2}\|\check{\mathcal{P}}_{Z}^{\perp}\hat{Z}\check{\mathcal{P}}_{Z}^{\perp}\|_{*}.$
@@
    $\|\hat{Z}-Z\|_{F}^{2}$ $+\left(\rho-\left\|E\right\|_{2}\right)\|\check{\mathcal{P}}_{Z}^{\perp}\hat{Z}\check{\mathcal{P}}_{Z}^{\perp}\|_{*}\leq\sqrt{d}\left(\rho+3\left\|E\right\|_{2}\right)\|\hat{Z}-Z\|_{F}.$
@@
    $\left|v^{\top}Zv\right|=\left|v^{\top}VV^{\top}ZVV^{\top}v\right|\leq\left\|Z\right\|_{2}\left\|V^{\top}v\right\|_{2}^{2}.$ (57)
@@
    Thus at most first $d$ eigenvalues of $\hat{Z}$ survive the soft-thresholding step
@@
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}(\hat{Z}_{\ell}-Z_{\ell})\Bigr\|_{2}\leq 11(\gamma\theta\vee\rho)\Bigl[{1\over m}\vee s\vee\theta\Bigr]^{1/2}.$ (59)
@@
Term I. We start by bounding the cosine similarity between perturbed eigenspaces in terms of the similarity of the true eigenspaces. By (33) and the triangular inequality, we have
$\left\|\tilde{V}_{\ell_{1}}^{\top}\tilde{V}_{\ell_{2}}\right\|_{2}$ $=\inf_{O_{1}\in\mathcal{O}_{d_{\ell_{1}}},\ O_{2}\in\mathcal{O}_{d_{\ell_{2}}}}\left\|\tilde{V}_{\ell_{1}}^{\top}\tilde{V}_{\ell_{2}}-\tilde{V}_{\ell_{1}}^{\top}V_{\ell_{2}}O_{2}+\tilde{V}_{\ell_{1}}^{\top}V_{\ell_{2}}O_{2}-O_{1}V_{\ell_{1}}^{\top}V_{\ell_{2}}O_{2}+O_{1}V_{\ell_{1}}^{\top}V_{\ell_{2}}O_{2}\right\|_{2}$ (60)
$\leq\inf_{O_{2}\in\mathcal{O}_{d_{\ell_{2}}}}\left\|\tilde{V}_{\ell_{2}}-V_{\ell_{2}}O_{2}\right\|_{2}+\inf_{O_{1}\in\mathcal{O}_{d_{\ell_{1}}}}\left\|\tilde{V}_{\ell_{1}}-V_{\ell_{1}}O_{1}\right\|_{2}+\left\|V_{\ell_{1}}^{\top}V_{\ell_{2}}\right\|_{2}$
$\leq 2\sqrt{2}\theta+s.$
By property 3 in Lemma B.2 and the threshold choice in (58), we have $\|\hat{Z}_{\ell}\|_{2}\leq\gamma$. Thus, by (32) and submultiplicativity
$\left\|\left(\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}-V_{\ell}V_{\ell}^{\top}\right)\hat{Z}_{\ell}\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}\right\|_{2}$ $\leq\|\tilde{V}_{\ell}\tilde{V}_{\ell}^{\top}-V_{\ell}V_{\ell}^{\top}\|_{2}\|\hat{Z}_{\ell}\|_{2}\leq 2\theta\gamma.$
@@
Therefore, by Lemma 4 in MacDonald et al. (2022)
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}I_{\ell}\Bigr\|_{2}\leq 2m^{-1/2}\theta\gamma\Bigl[1+m(2\sqrt{2}\theta+s)\Bigr]^{1/2}\leq 5\theta\gamma\Bigl[{1\over m}\vee\theta\vee s\Bigr]^{1/2},$
where we used $(2+2\sqrt{2})^{1/2}<2.5$.
Term II. With the choice of the thresholds in (B.2), we have by property 1 in Lemma B.2 that $\|\Delta_{\ell}\|_{2}\leq 2\rho$
$\|V_{\ell_{1}}V_{\ell_{1}}^{\top}\Delta_{\ell_{1}}V_{\ell_{1}}V_{\ell_{1}}^{\top}\cdot V_{\ell_{2}}V_{\ell_{2}}^{\top}\Delta_{\ell_{2}}V_{\ell_{2}}V_{\ell_{2}}^{\top}\|_{2}\leq\|\Delta_{\ell_{1}}\|_{2}\|V_{\ell_{1}}^{\top}V_{\ell_{2}}\|_{2}\|\Delta_{\ell_{2}}\|_{2}\leq 4s\rho^{2}.$
Thus, by Lemma 4 in MacDonald et al. (2022),
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}II_{\ell}\Bigr\|_{2}\leq 2m^{-1/2}\rho\Bigl[1+ms\Bigr]^{1/2}\leq 3\rho\Bigl[{1\over m}\vee s\Bigr]^{1/2}.$
Term III. Similarly to Term II, we can establish by substituting $\|\tilde{V}_{\ell_{1}}\tilde{V}_{\ell_{2}}\|_{2}$ with $\|V_{\ell_{1}}V_{\ell_{2}}\|_{2}$ in (61):
$\Bigl\|{1\over m}\sum_{\ell=1}^{m}III_{\ell}\Bigr\|_{2}\leq 2m^{-1/2}\theta\gamma\Bigl[1+ms\Bigr]^{1/2}\leq 3\theta\gamma\Bigl[{1\over m}\vee s\Bigr]^{1/2}.$
Lemma B.4 (Hard thresholding with noise).
In notation of Lemma B.2 with $\hat{Z}:=[\tilde{Z}]_{d}$
  1. 1.
    The spectral norm of the difference can be bounded as
    $\|\hat{Z}-Z\|_{2}\leq 2\|E\|_{2}$
  2. 2.
    If additionally $|\gamma_{d}(Z)|\geq 4\|E\|_{2}$, it holds
    $\|\hat{Z}-Z\|_{2}\leq{19\|Z\|_{2}\over|\gamma_{d}(Z)|}\|\mathcal{P}_{V}E\|_{2}.$
Proof.
  1. 1.
    Consider the eigendecomposition $Z+E=\tilde{U}\tilde{\Gamma}\tilde{U}^{\top}$. Then, the eigendecomposition of $\hat{Z}$ can be written as $\hat{Z}=\tilde{U}[\tilde{\Gamma}]_{d}\tilde{U}^{\top}$ with
    $\hat{Z}-Z=\tilde{U}[\tilde{\Gamma}]_{d}\tilde{U}^{\top}-\left(\tilde{U}\tilde{\Gamma}\tilde{U}^{\top}-E\right)=\tilde{U}\left([\tilde{\Gamma}]_{d}-\tilde{\Gamma}\right)\tilde{U}^{\top}+E$
    Therefore, by Weil’s inequality
    $\|\hat{Z}-Z\|_{2}\leq|\gamma_{d+1}(Z+E)|+\|E\|_{2}\leq|\gamma_{d+1}(Z)|+2\|E\|_{2}=2\|E\|_{2}$
  2. 2.
    Decompose the error as follows:
    $\hat{Z}-Z$ $=\mathcal{P}_{\tilde{V}}\tilde{Z}\mathcal{P}_{\tilde{V}}-\mathcal{P}_{V}Z\mathcal{P}_{V}=(\mathcal{P}_{\tilde{V}}-\mathcal{P}_{{V}})\tilde{Z}\mathcal{P}_{\tilde{V}}+\mathcal{P}_{V}E\mathcal{P}_{\tilde{V}}+\mathcal{P}_{V}Z(\mathcal{P}_{\tilde{V}}-\mathcal{P}_{V})$
    By Corollary B.1 and (32),
    $\|\mathcal{P}_{V}-\mathcal{P}_{\tilde{V}}\|_{2}=\|\tilde{V}\tilde{V}^{\top}-V^{\top}V^{\top}\|_{2}\leq 2\sin\Theta(V,\tilde{V})\leq{2\|\mathcal{P}_{V}E\|_{2}\over|\gamma_{d}(Z)|-3\|E\|_{2}}$
    Notice also that $\|\tilde{Z}\|_{2}\leq\|Z\|_{2}+\|E\|_{2}\leq\|Z\|_{2}+|\gamma_{d}(Z)|/4=5\|Z\|_{2}/4$. Therefore, by triangular inequality and submultiplicativity:
    $\|\hat{Z}-Z\|_{2}$ $\leq\|\mathcal{P}_{V}E\|_{2}+(\|Z\|_{2}+\|\tilde{Z}\|_{2}){2\|\mathcal{P}_{V}E\|_{2}\over|\gamma_{d}(Z)|-3\|E\|_{2}}$
    $\leq\|\mathcal{P}_{V}E\|_{2}+{9\|Z\|_{2}\over 4}{2\|\mathcal{P}_{V}E\|_{2}\over|\gamma_{d}(Z)|/4}$
    $\leq{19\|Z\|_{2}\over|\gamma_{d}(Z)|}\|\mathcal{P}_{V}E\|_{2}.$